The steel rail road track is 20 cm long at 10 degree celsius. How much longer is the track at 20 degree celsius? The coefficient of linear expansion of steel is 1.27x10-5 -K.

1
Expert's answer
2022-04-11T12:06:16-0400
"\Delta l=l_0\alpha \Delta T"

"\Delta l=20\:{\rm cm}*1.27*10^{-5}*(20-10)\\
=2.54*10^{-3}\: \rm cm"